27. Toluene, C6H5CH3, is given below. Draw the maj

27. Toluene, C6H5CH3, is given below. Draw the major organic product for each reaction. Assume 1:1 reaction stoichiometry. If more than one product is possible, only one structure needs to be drawn.
Answer #1
Concepts and reason

This problem is based on the concept of reactions of toluene.

Toluene is an aromatic compound which contains a benzene ring and has a methyl group attached to this benzene ring. This methyl group increase the electron density of benzene ring due to its electron donating tendency through inductive effect.

Fundamentals

Toluene on reaction with potassium permanganate (KMnO4)\left( {{\rm{KMn}}{{\rm{O}}_4}} \right) in acidic condition gets oxidized to the benzoic acid which is a meta directing compound. Toluene on reaction with

Bromine in the presence of ferric bromide (FeBr3)\left( {{\rm{FeB}}{{\rm{r}}_3}} \right) gives a brominated product in which bromine atom is attached to the para position to the methyl group present in the toluene.

[Part a]

The reaction of toluene with potassium permanganate (KMnO4)\left( {{\rm{KMn}}{{\rm{O}}_4}} \right) is shown below:

COOH
1.) KMnO,,H,O, A

The complete reaction sequence is given below:

COOH
1.) KMnO,,H,O, A
2.) Brą, FeBrz

[Part b]

The reaction of toluene with bromine in the presence of ferric chloride is shown below:

1.) Bry, FeBr,
لا
Br

The complete reaction sequence is given below:

COOH
1.) Bry, FeBr
2.) KMnO,H,O, A

Ans: Part a

The structure of the product is as follows:

COOH

Part b

The structure of the product is as follows: